We are currently monitoring the situation and awaiting the final report regarding the power outage caused by PDU#3 (approx. 1/3 of the power capacity in the data center MUC1 was affected).
13:00 Power has been fully restored, we are now watching the boot procedure of the servers to come online again. Not all services are restored yet, but will be soon.
12:43 All switches are operational again, our technicians are trying to switch over servers without power to the two power circuits still working in the data center without causing an overload.
12:20 One third of the power is offline in the data center MUC1 (Level3) caused by a defect PDU. Our technicians are already in the data center trying to get all switches back online so that at least customers with redundant power supplies are online again. We still have no ETA from the data center operator, but will update the status page as soon as possible.
11:44 We have detected a partial power outage in the data center #1.